# Jack Hong   , Chenxiao Zhao, ChengLIn Zhu, Weiheng Lu, Guohai Xu   , XingYu

### v2

@@ 5.1 Implementation Details
We conduct training in two stages: cold start SFT and reinforcement learning. The backbone model is Qwen2.5-VL-7B (bai2025qwen2). For SFT, we train with a batch size of 128 and a learning rate of $1\times 10^{-5}$. Model is optimized for 3 epochs using AdamW (loshchilov2017decoupled) optimizer with cosine learning rate decay. For RL, we adopt DAPO (yu2025dapo) as the optimization algorithm, with a batch size of 256 and 16 rollouts per prompt. The KL coefficient is set to $0.0$, and the maximum response length is capped at $16,384$ tokens. The learning rate is $1\times 10^{-6}$, and the upper and lower clip ratios are $0.30$ and $0.20$, respectively. We utilize VLMEvalKit (duan2024vlmevalkit) to conduct all the evaluation, except for RealX-Bench, so the performance of DeepEyes may be a little different from (zheng2025deepeyes).
@@ 5.3 Results on Other Benchmarks
Real-World & OCR & Chart Understanding. We evaluate DeepEyesV2 across three categories of benchmarks: real-world understanding, OCR, and chart understanding. For comparison, we include two kinds of models: (i) open-source general-purpose MLLMs, including LLaVA-OneVision (li2024llava), Qwen2.5-VL (bai2025qwen2), and InternVL3 (zhu2025internvl3); and (ii) grounded reasoning models, such as DeepEyes (zheng2025deepeyes) and Thyme (zhang2025thyme). DeepEyes performs fine-grained perception by cropping the target region, while Thyme manipulates images through executable code. Compared to base model Qwen2.5-VL-7B, DeepEyesV2 demonstrates substantial performance gains, and even surpasses Qwen2.5-VL-32B in some benchmarks (Table 4), highlighting the effectiveness of tool-augmented reasoning. Moreover, it consistently outperforms existing grounded reasoning models. These results indicate that dynamic tool invocation enables model to extract fine-grained details, thereby improving real-world scene comprehension.
Multimodal Reasoning. We further evaluate DeepEyesV2 on mathematical reasoning benchmarks to assess its strong reasoning capability. As shown in Table 4, we compare DeepEyesV2 against existing open-source MLLMs, such as Qwen2.5-VL (bai2025qwen2), text-only multimodal reasoning models, including MM-Eureka (meng2025mm), and grounded reasoning models, such as DeepEyes (zheng2025deepeyes) and Thyme (zhang2025thyme). DeepEyesV2 consistently outperforms these alternatives, and notably achieves stronger results than text-only multimodal reasoning models, underscoring the benefit of tool use for enhancing mathematical reasoning.
Online Searching. To further examine the effectiveness of external information acquisition, we evaluate DeepEyesV2 on search-oriented benchmarks. These datasets encompass knowledge-intensive visual question answering, fact verification, and multimodal retrieval-based reasoning, all of which require models to go beyond perceptual understanding and actively retrieve external evidence. For comparison, we benchmark DeepEyesV2 against both general-purpose MLLMs such as Qwen2.5-VL, Gemini 2.5 Pro, and GPT4o, as well as models where search capability is incorporated (jiang2024mmsearch; geng2025webwatcher). As shown in Table 4, DeepEyesV2 demonstrates superior search capabilities, achieving consistently higher accuracy across all benchmarks.
@@ A.1 Training Data
For perception-oriented tasks, we include V* [wu2024v], ArxivQA [li2024multimodal], Pixmo Counting [deitke2024molmo], TallyQA [acharya2019tallyqa], and SeekWorld [seekworld2025], covering a wide range of scenarios such as object recognition, visual counting, and chart interpretation. For reasoning tasks, we adopt ReVisual [chen2025advancing] to provide complex reasoning problems, and additionally incorporate MathCoder [wang2023mathcoder] and Retool [feng2025retool] to supplement with executable code-based problem-solving examples. Besides, we also include long Chain-of-Cot (CoT) reasoning data in cold start subset. For search-related tasks, we employ MMSearch-R1 [wu2025mmsearch], which includes both image-based and text-based retrieval questions. We further include data from VGR [wang2025vgr], Chain-of-Focus [zhang2025chain], and VLM-R3 [jiang2025vlm] to strengthen the reinforcement learning corpus.
@@ A.2 Evaluation Protocol
Real-World & OCR & Chart Understanding. For real-world understanding, we adopt V* [wu2024v], HRBench [wang2025divide], MME-RealWorld [zhang2024mme], and TreeBench [wang2025traceable]; for OCR, we use OCRBench [liu2024ocrbench] and Seed-Bench-2-Plus [li2024seed]; and for chart reasoning, we evaluate on CharXiv [wang2024charxiv] and ChartQA [masry2022chartqa]. For comparison, we include two kinds of models: (i) open-source general-purpose MLLMs, including LLaVA-OneVision [li2024llava], Qwen2.5-VL [bai2025qwen2], and InternVL3 [zhu2025internvl3]; and (ii) grounded reasoning models, such as Pixel-Reasoner [su2025pixel], DeepEyes [zheng2025deepeyes] and Thyme [zhang2025thyme].
@@
Multimodal Reasoning. We include MathVista [lu2023mathvista], MathVerse [zhang2024mathverse], MathVision [wang2024measuring], WeMath [qiao2024we], and LogicVista [xiao2024logicvista]. We compare DeepEyesV2 against existing open-source MLLMs, such as Qwen2.5-VL [bai2025qwen2], text-only multimodal reasoning models, including MM-Eureka [meng2025mm], ThinkLite [wang2025sota], VL-Rethinker [wang2025vl], and VLAA-Thinker [chen2025sft], and grounded reasoning models, such as DeepEyes [zheng2025deepeyes] and Thyme [zhang2025thyme]
@@
Online Searching. We compare DeepEyesV2 on FVQA-test [wu2025mmsearch], InfoSeek [chen2023can], MMSearch [jiang2024mmsearch], and SimpleVQA [cheng2025simplevqa]. We benchmark DeepEyesV2 against both general-purpose MLLMs such as Qwen2.5-VL [bai2025qwen2], Gemini 2.5 Pro [comanici2025gemini], and GPT4o [hurst2024gpt], as well as models where search capability is incorporated [jiang2024mmsearch, geng2025webwatcher].
A.3 Tool Taxonomy
@@ A.5 Prompt
A.6 More Cases
We show more cases in Figure11,12.

### v3

@@ 5.1 Implementation Details
We conduct training in two stages: cold start SFT and reinforcement learning. The backbone model is Qwen2.5-VL-7B (Bai et al., 2025). For SFT, we train with a batch size of 128 and a learning rate of $1\times 10^{-5}$. Model is optimized for 3 epochs using AdamW (Loshchilov and Hutter, 2017) optimizer with cosine learning rate decay. For RL, we adopt DAPO (Yu et al., 2025) as the optimization algorithm, with a batch size of 256 and 16 rollouts per prompt. The KL coefficient is set to $0.0$, and the maximum response length is capped at $16,384$ tokens. The learning rate is $1\times 10^{-6}$, and the upper and lower clip ratios are $0.30$ and $0.20$, respectively. We utilize VLMEvalKit (Duan et al., 2024) to conduct all the evaluation, except for RealX-Bench, so the performance of DeepEyes may be a little different from (Zheng et al., 2025b).
@@ 5.3 Results on Other Benchmarks
Real-World & OCR & Chart Understanding. We evaluate DeepEyesV2 across three categories of benchmarks: real-world understanding, OCR, and chart understanding. For comparison, we include two kinds of models: (i) open-source general-purpose MLLMs, including LLaVA-OneVision (Li et al., 2024a), Qwen2.5-VL (Bai et al., 2025), and InternVL3 (Zhu et al., 2025); and (ii) grounded reasoning models, such as DeepEyes (Zheng et al., 2025b) and Thyme (Zhang et al., 2025b). DeepEyes performs fine-grained perception by cropping the target region, while Thyme manipulates images through executable code. Compared to base model Qwen2.5-VL-7B, DeepEyesV2 demonstrates substantial performance gains, and even surpasses Qwen2.5-VL-32B in some benchmarks (Table 4), highlighting the effectiveness of tool-augmented reasoning. Moreover, it consistently outperforms existing grounded reasoning models. These results indicate that dynamic tool invocation enables model to extract fine-grained details, thereby improving real-world scene comprehension.
Multimodal Reasoning. We further evaluate DeepEyesV2 on mathematical reasoning benchmarks to assess its strong reasoning capability. As shown in Table 4, we compare DeepEyesV2 against existing open-source MLLMs, such as Qwen2.5-VL (Bai et al., 2025), text-only multimodal reasoning models, including MM-Eureka (Meng et al., 2025), and grounded reasoning models, such as DeepEyes (Zheng et al., 2025b) and Thyme (Zhang et al., 2025b). DeepEyesV2 consistently outperforms these alternatives, and notably achieves stronger results than text-only multimodal reasoning models, underscoring the benefit of tool use for enhancing mathematical reasoning.
Online Searching. To further examine the effectiveness of external information acquisition, we evaluate DeepEyesV2 on search-oriented benchmarks. These datasets encompass knowledge-intensive visual question answering, fact verification, and multimodal retrieval-based reasoning, all of which require models to go beyond perceptual understanding and actively retrieve external evidence. For comparison, we benchmark DeepEyesV2 against both general-purpose MLLMs such as Qwen2.5-VL, Gemini 2.5 Pro, and GPT4o, as well as models where search capability is incorporated (Jiang et al., 2024; Geng et al., 2025). As shown in Table 4, DeepEyesV2 demonstrates superior search capabilities, achieving consistently higher accuracy across all benchmarks.
@@ A.1 Training Data
For perception-oriented tasks, we include V* [54], ArxivQA [27], Pixmo Counting [11], TallyQA [1], and SeekWorld [43], covering a wide range of scenarios such as object recognition, visual counting, and chart interpretation. For reasoning tasks, we adopt ReVisual [6] to provide complex reasoning problems, and additionally incorporate MathCoder [48] and Retool [14] to supplement with executable code-based problem-solving examples. Besides, we also include long Chain-of-Cot (CoT) reasoning data in cold start subset. For search-related tasks, we employ MMSearch-R1 [53], which includes both image-based and text-based retrieval questions. We further include data from VGR [46], Chain-of-Focus [58], and VLM-R3 [20] to strengthen the reinforcement learning corpus.
@@ A.2 Evaluation Protocol
Real-World & OCR & Chart Understanding. For real-world understanding, we adopt V* [54], HRBench [50], MME-RealWorld [60], and TreeBench [44]; for OCR, we use OCRBench [33] and Seed-Bench-2-Plus [25]; and for chart reasoning, we evaluate on CharXiv [52] and ChartQA [37]. For comparison, we include two kinds of models: (i) open-source general-purpose MLLMs, including LLaVA-OneVision [24], Qwen2.5-VL [4], and InternVL3 [65]; and (ii) grounded reasoning models, such as Pixel-Reasoner [42], DeepEyes [64] and Thyme [59].
@@
Multimodal Reasoning. We include MathVista [36], MathVerse [57], MathVision [47], WeMath [40], and LogicVista [55]. We compare DeepEyesV2 against existing open-source MLLMs, such as Qwen2.5-VL [4], text-only multimodal reasoning models, including MM-Eureka [38], ThinkLite [51], VL-Rethinker [45], and VLAA-Thinker [5], and grounded reasoning models, such as DeepEyes [64] and Thyme [59]
@@
Online Searching. We compare DeepEyesV2 on FVQA-test [53], InfoSeek [7], MMSearch [21], and SimpleVQA [9]. We benchmark DeepEyesV2 against both general-purpose MLLMs such as Qwen2.5-VL [4], Gemini 2.5 Pro [10], and GPT4o [18], as well as models where search capability is incorporated [21, 16].
Model Tool Param Size Real-World Understanding OCR Chart V* Bench HRBench 4K HRBench 8K MME- RealWorld Tree Bench OCR Bench SEED 2 Plus CharXiv descriptive CharXiv reasoning Chart QA Open-source Models LLaVA-OV 7B 75.4 63.0 59.8 57.4 37.3 - - - - 80.0 Qwen2.5-VL 7B 78.5 71.6 67.9 57.3 37.0 864 70.4 72.7 40.2 86.2 Qwen2.5-VL 32B 80.6 74.1 69.9 - 42.5 - 72.4 83.2 48.0 - InternVL3 8B 81.2 70.0 69.3 - 38.8 880 69.7 73.6 37.6 86.6 Proprietary Models with Tools GPT-4o Code - 58.6 60.6 55.1 - 48.1 822 71.9 85.9 47.5 - Gemini 2.5 Pro Code - 79.6 86.9 - 71.6 49.1 881 75.0 92.6 67.8 - Grounded Reasoning Models Pixel-Reasoner Crop 7B 84.3 74.0 66.9 64.4 39.0 - - - - - DeepEyes Crop 7B 85.6 75.1 72.6 - 37.5 - - - - - Thyme Code 7B 82.2 77.0 72.0 64.8 - 863 - - - 86.1 Agentic Multimodal Model DeepEyesV2 General 7B 81.8 77.9 73.8 64.9 42.5 882 70.5 78.6 48.9 88.4 $\Delta$ (vs Qwen2.5-VL-7B) +3.3 +6.3 +5.9 +7.6 +5.5 +18 +0.1 +5.9 +8.7 +2.2
Table 8: Results on multimodal reasoning benchmarks.
Model Tool Param Size MathVista MathVerse MathVision WeMath DynaMath LogicVista Open-source Models LLaVA-OV 7B 58.6 19.3 18.3 20.9 - 33.3 Qwen-2.5-VL 7B 68.3 45.6 25.6 34.6 53.3 45.9 InternVL3 8B 71.6 39.8 29.3 37.1 - 44.1 Proprietary Models with Tools GPT-4o Code - 59.3 54.0 - 41.6 61.9 51.2 Gemini 2.5 Pro Code - 83.0 81.4 - - - - Text-only Reasoning Models MM-Eureka 7B 72.6 - 28.1 21.8 - 46.3 ThinkLite 7B 71.6 - 24.6 41.8 - 42.7 VL-Rethinker 7B 73.7 - 28.4 36.3 - 42.7 VLAA-Thinker 7B 71.7 - 24.2 35.7 - 45.9 Grounded Reasoning Models DeepEyes Crop 7B 70.1 47.3 26.6 38.9 55.0 47.7 Thyme Code 7B 70.0 - 27.6 39.3 - 49.0 Agentic Multimodal Model DeepEyesV2 General 7B 71.9 52.7 28.9 38.1 57.2 48.7 $\Delta$ (vs Qwen2.5-VL-7B) +3.6 +7.1 +3.3 +3.5 +3.9 +2.8
A.3 Performance Comparison with Proprietary Models
In Table 7 and  8, we compare DeepEyesV2 with existing proprietary models by using the same prompt as DeepEyesV2, DeepEyesV2 achieves performance comparable to GPT-4o.
Figure 10: Tool Usage Pattern during RL.
A.4 Tool Usage Pattern Evolution
To further understand how tool usage patterns evolve during RL training, Figure 10 presents category-wise statistics of code tools across training stages. Clear trends emerge: (i) the proportion of “Mark” tools remains largely stable, with minimal fluctuation; (ii) the proportion of “Crop” tools rises initially and then declines as training progresses; and (iii) the “Numerical analysis” and “Other” categories exhibit the opposite trajectory—decreasing at first and subsequently increasing. These trends in tool usage are broadly consistent with the changes in tool usage shown in Figure 8.
Qwen2.5-VL 7B DeepEyes DeepEyesV2-SFT DeepEyesV2-RL 16.0 17.3 18.7 20.8
Table 9: Zero-shot results on TIR-Bench
DeepEyesV2 DeepEyesV2 w/ Rotate 33.5 38.3
Table 10: Zero-shot results on OCR subset of TIR-Bench with addition tools
Figure 11: Zero-shot tool utilization example of DeepEyesV2 on TIR-Bench.
A.5 Zero-Shot Generalization of Tool Usage
We conduct zero-shot testing of DeepEyesV2 on TIR-Bench [28]. The tasks and required tool categories in TIR-Bench are not included in our cold-start and RL training data, therefore TIR-Bench can serve as an excellent platform for zero-shot testing of DeepEyesV2 to illustrate its generalization capability in tool invocation.
Performance Comparison. The performance of DeepEyesV2 on TIR-Bench is summarized in Table 10. Specifically, DeepEyesV2-SFT represents the model after the cold-start stage, and DeepEyesV2-RL represents the model after RL training. It is evident that DeepEyesV2 substantially outperforms the baseline (Qwen2.5-VL 7B) thanks to tool integration. This provides compelling evidence of DeepEyesV2’s strong generalization ability, confirming that great performance improvements persist even when applied to unseen tools and tasks.
New Tool Visualization. We present examples of zero-shot tool utilization from TIR-Bench in Figure 11, comparing model performance before and after RL across tasks such as OCR, image rotation, and maze solving. We omit the specific options for each question and only display the core code segments. Prior to RL, lacking exposure to relevant tasks, model merely represent the input image. Conversely, the post-RL model demonstrates robust generalization capabilities. Specifically, for OCR task, DeepEyesV2 employs grayscaling and dilation to enhance character clarity, followed by cropping individual digits for recognition. For the rotation task, DeepEyesV2 rotates the image to determine its original orientation angle. Notably, these tasks and tools are absent from both our cold-start and RL datasets; yet, without additional training, DeepEyesV2 successfully comprehends and utilizes these novel tools. Furthermore, we highlight an interesting maze scenario where DeepEyesV2 generates code to simulate pathfinding, thereby verifying each option. Although the code contains minor imperfections, we think this strongly evidences DeepEyesV2’s generalization potential with new tools. This comparison of tool usage pre- and post-RL effectively highlights DeepEyesV2’s exceptional adaptability to unseen tasks and tools.
Function Call Generalization. Furthermore, beyond the code generalization capabilities, DeepEyesV2 demonstrates strong generalization in function calling. To evaluate this, we equip DeepEyesV2 with a rotation tool on a subset of the Rotated OCR task from TIR-Bench, requiring model to utilize the tool via function calls rather than by writing code. Since images in this task are rotated, model must perform rotation prior to OCR recognition. It is worth noting that in addition to the rotation function call, DeepEyesV2 retains the ability to invoke other tools via code generation. As shown in Table 10, providing the rotation function call leads to further performance improvements, with DeepEyesV2 invoking the rotation tool in $15\%$ of instances. This demonstrates DeepEyesV2’s strong generalization capability regarding newly added tools.
A.6 Tool Taxonomy
@@ A.8 Prompt
A.9 More Cases
We show more cases in Figure13,14.